If initial velocity is 64 ft/s what is the maximum height...?
A soccer ball is kicked off the ground at 64 ft/s... What is the max height?
Use projectile function h(t)=-16t^2+Vt+H
H=initial height
V=initial velocity
t=time
h=height

Answer by stanbon(75887)

If initial velocity is 64 ft/s what is the maximum height...?
A soccer ball is kicked off the ground at 64 ft/s... What is the max height?
Use projectile function h(t)=-16t^2+Vt+H
------
Height at time = t is h(t) = -16t^2 + 64t + 0
---------------
Max height occurs when t = -b/(2a) = -64/(2*-16) = 2 seconds
---
Max height = h(2) = -16*2^2 + 64*2 + 0 = -64+2*64 = 64 ft
-----------------------
Cheers,
Stan H.
------------------
H=initial height
V=initial velocity
t=time
h=height